Page 1 Question 1 Suppose we take repeated random samples of size 20 from a population with a Select all that apply. mean ofQuestion 3 Suppose that a candy company makes a candy Select one bar whose weight is supposed to be 50 grams, answer. but inQuestion 4 In which of the following scenarios would the Select all distribution of the sample mean x-bar be that apply. normQuestion 6 The number of hours a light bulb burns before Select one failing varies from bulb to bulb. The answer. distributioQuestion 7 A machine is designed to fill 16-ounce bottles Type of shampoo. When the machine is working numbers in properly, tQuestion 8 Suppose that the distribution for total Type amounts spent by students vacationing for a numbers in week in FloridQuestion 2 Pictured below (in scrambled order) are three Select one histograms. One of them represents a answer. population d
Answer #1

Solution:-

1. option A. The distribution will be normal as long as the population distribution is normal.
option C. The distribution's mean is the same as the population mean 60.

3. option D.it is extreamly unlike for this to occur;the probability is very close to 0.
Explanation : P(diff > 4) = P(Z > 4) = P(Z>4) =0.0000

4.option B. We take repeated random samples of size 15 from a population that is normally distributed.
option C. We take repeated random samples of size 50 from a population of unknown shape.

5. P(X < 4.2) = P(Z < (4.2-4)/(1.1/sqrt(100)) = 0.9656